# A Little Depth Goes a Long Way: The Expressive Power of Log-Depth Transformers

### v2

@@ 1 Introduction
A line of recent work analyzing the intrinsic computational power of transformers, the neural architecture behind today’s immensely successful large language models (LLMs), has established that, with fixed depth, transformers cannot express many simple problems outside the complexity class $\mathsf{TC}^{0}$, including recognizing regular languages and resolving connectivity of nodes in a graph (Merrill and Sabharwal, 2023a; Chiang et al., 2023). These problems conceivably underlie many natural forms of sequential reasoning, such as state tracking (Liu et al., 2023; Merrill et al., 2024) and resolving logical inferences across long chains (Wei et al., 2022). Thus, these results suggest inherent limitations on the types of reasoning transformer classifiers can perform. Yet, these findings come with an important caveat: even if transformers cannot solve such problems exactly for inputs of arbitrary lengths, they may still be able to solve them over inputs up to some bounded length. This perspective, coupled with the fact that treating depth as fixed is crucial to prior analyses placing transformers in $\mathsf{TC}^{0}$, motivates two related questions about depth as an important resource for a transformer, in relation to the context length over which it reasons:
@@
We address these questions by analyzing the expressive power of “universal” transformers (also called “looped” transformers) where a fixed model is given dynamic depth by repeating a block of middle layers a variable number of times (Dehghani et al., 2019; Yang et al., 2024; Geiping et al., 2025). We capture the regime where depth grows minimally with context length, with the middle layers repeated $\Theta(\log n)$ times on contexts of length $n$. We prove that even such highly uniform transformers, when allowed log-depth, can recognize regular languages [^1] While Liu et al. (2023) provided a similar result, their construction relied heavily on an unconventional, non-uniform transformer architecture, requiring a different set of model weights for each input length. In contrast, our result holds in a stronger setting of highly uniform transformers—where model weights are not only fixed (independent of input length) as in practice but even shared across blocks of layers, enabling effective inference-time scaling. Our formal model incorporates standard architectural choices like residual connection and layer norm. We supplement our stronger theoretical findings with matching empirical learnability results. and solve graph connectivity, two important reasoning problems known to be beyond fixed-depth transformers (Merrill and Sabharwal, 2023a). Our core technical contribution enabling this is Lemma 1, that fully uniform transformers can compute division and remainder of small integers. This not only obviates the need for non-uniformity and special positional encodings relied upon in prior work, it is also an interesting finding on its own.
@@
Figure 1: To recognize a regular language over inputs of length $n$, the depth of a universal transformer can grow $\Theta(\log n)$ by Theorem 1. On the other hand, width must grow superpolynomially (Theorem 3), and the number of chain-of-thought steps must be superlogarithmic (Theorem 4). The precise depth and width coefficients plotted here were obtained experimentally in Section 7.
@@
Our result has two interesting interpretations. First, it directly shows that, by dynamically increasing their depth as $\Theta(\log n)$ on inputs of length $n$, one can construct transformers to solve regular language recognition (Theorem 1) and graph connectivity (Theorem 2) for arbitrary context length. In contrast, chain-of-thought (CoT) steps, used for additional test-time compute by newest LLMs such as OpenAI o1 (OpenAI, 2024) and DeepSeek-R1 (DeepSeek AI, 2025), must be scaled superlogarithmically (Theorem 4) to solve these problems, and width must be scaled superpolynomially (Theorem 3), as shown in Figure 1. Thus scaling depth more efficiently allows solving these reasoning problems compared to scaling width or using CoT.
@@
Second, a universal transformer unrolled to a fixed (independent of input size) depth $d$ is a special case of a standard $d$-depth transformer, namely one with a highly uniform structure (parameters are shared across layers). Thus, our result shows that standard transformers with a fixed depth $d$ can recognize regular languages (Corollary 1.2) and solve graph connectivity problems (Corollary 2.1) as long as one cares only about bounded inputs of size $2^{O(d)}$. This allows us to quantify how many layers are necessary for a desired input size. For instance, it follows from Corollaries 1.2 and 2.1 that with a depth of only $32$, such as in smaller models like LLaMA 3.1 7B (Meta AI, 2024) and OLMo 7B (Ai2, 2024), transformers can recognize regular languages up to strings of length $107$ and solve connectivity for graphs with up to $128$ vertices. A depth of $80$ (such as in LLaMA 3.1 70B) makes these input size limits practically unbounded: strings of length up to $440$K and graphs with up to $2.1$B vertices, respectively. Empirically, our experiments supplementing our theoretical findings demonstrate that scaling depth as $\Theta(\log n)$ is necessary and sufficient for learning to recognize hard regular languages.
@@ 2 Preliminaries: Universal Transformers
We consider $(s,r,t)$-universal transformers, which are defined to have $s$ fixed initial layers at the start, a sequence of $r$ layers that is repeated some number of times based on the input length, and a sequence of $t$ fixed final/terminal layers. Thus, an $(s,r,t)$-universal transformer unrolled $d(n)$ times for input length $n$ has a total of $s+rd(n)+t$ layers. Geiping et al. (2025) empirically explored such transformers for scaling test-time computation for reasoning problems. A standard $d$-layer transformer is $(d,0,0)$-universal (equivalently, $(0,0,d)$-universal), while a standard universal transformer (Dehghani et al., 2019; Yang et al., 2024) is $(0,1,0)$-universal.
@@
    An embedding projection matrix $\mathbf{E}:\Sigma\to\mathbb{Q}^{m}$ and positional encoding function $\pi:\mathbb{N}\to\mathbb{Q}^{m}$, which we assume separates $1$ from other indices (Merrill and Sabharwal, 2024); [^2] We use rationals $\mathbb{Q}$ instead of $\mathbb{R}$ so that the model has a finite description. All our simulations go through as long as at least $c\log n$ bits are used to represent rationals, similar in spirit to log-precision floats used in earlier analysis (Merrill and Sabharwal, 2023a, b).
@@
We next define how the transformer maps a sequence $w_{1}\cdots w_{n}\in\Sigma^{n}$ to an output value $y\in\Sigma$; to do so, we will always specify that the transformer is unrolled to a specific depth function $d(n)$, which we will consider to be $d(n)=\lceil\log n\rceil$. [^3] Following computer science conventions, $\log n\triangleq\log_{2}n$. The computation is inductively defined by the residual stream $\mathbf{h}_{i}$: a cumulative sum of all layer outputs at each token $i$. In the base case, the residual stream $\mathbf{h}_{i}$ is initialized to $\mathbf{h}^{0}_{i}=\mathbf{E}(w_{i})+\pi(i)$. We then iteratively compute $s+rd(n)+t$ more layers, deciding which layer to use at each step as follows:
$L^{\ell}=\begin{cases}s\textrm{-layer}\;\ell&\textrm{if}\;1<\ell\leq s\\ r\textrm{-layer}\;(\ell-s)\bmod r&\textrm{if}\;s<\ell\leq s+rd(n)\\ t\textrm{-layer}\;\ell-s-rd(n)&\textrm{otherwise.}\end{cases}$
@@
An $(s,r,t)$-transformer unrolled to some fixed depth can be viewed as a “uniform” special case of a fixed-depth transformer. Thus, constructions of dynamic-depth transformers (depth $d(n)$ for inputs of length $n$) imply that, given any bounded context length $N$, there also exists a fixed-depth transformer with depth $d(N)$ for the task at hand. The fact that this can be done with a looped transformer with dynamic depth is, in fact, a stronger condition that shows the construction is uniform, which is formally important as non-uniform models of computation can have very strong and unrealistic power (cf. Merrill et al., 2022). In this way, our results about looped transformers will provide insights about standard, non-looped transformers with bounded context lengths.
@@
To make Definition 1 well-defined, we will next describe the structure of the self-attention and feedforward sublayers that make up the structure of each transformer layer. Our definition of the transformer will have two minor differences from practice:
@@
    Masked pre-norm: We assume standard pre-norm (Xiong et al., 2020) but add a learned mask vector that can select specific dimensions of the residual stream for each layer’s input.
Each sublayer will take as input a sequence of normalized residual stream values $\mathbf{z}_{i}=\mathsf{layer\_norm}(\mathbf{m}\mathbf{h}_{i})$, where layer-norm can be standard layer-norm (Ba et al., 2016) or RMS norm (Zhang and Sennrich, 2019). The sublayer then maps $\mathbf{z}_{1},\ldots,\mathbf{z}_{n}$ to a sequence of updates to the residual stream $\delta_{1},\ldots,\delta_{n}$, and the residual stream is updated as $\mathbf{h}^{\prime}_{i}=\mathbf{h}_{i}+\delta_{i}$.
@@ 2.1 Memory Management in Universal Transformers
Appendix A discusses in detail how $\{-1,0,1\}$ values can be stored directly in the residual stream, while a general scalar $z$ can be stored either as $\psi(z)=\langle z,1,-z,-1\rangle$ in its unnormalized form or as the unit vector $\phi(z)=\psi(z)/\sqrt{z^{2}+1}$ in its normalized form (“layer-norm hash”; cf. Merrill and Sabharwal, 2024). Importantly, however $z$ is stored, when it is read using masked pre-norm, we obtain $\phi(z)$. In Appendix A, we show how numerical values represented using $\psi$ or $\phi$ can be easily written (Lemma 4), read (Lemma 2), and deleted (Lemmas 5 and 6) from the residual stream. We will leverage these operations heavily in our theoretical constructions.
@@ 3 Fixed Depth Transformers Can Divide Small Integers
Let $a_{i},b_{i},c_{i},m\in\mathbb{Z}^{\geq 0}$ be such that $a_{i}=b_{i}m+c_{i}$ where $a_{i}\leq i$ and $c_{i}<m$. Suppose $\psi(i)$, $\psi(m)$, and $\phi(a_{i})$ (or $\psi(a_{i})$) are present in the residual stream of a transformer at each token $i$. Then, there exists a 7-layer transformer with causally masked attention and masked pre-norm that, on any input sequence, adds $\phi(b_{i})$ and $\phi(c_{i})$ to the residual stream at each token $i$.
@@
The overall idea is as follows. In the first layer, each position $i$ outputs an indicator of whether it’s a multiple of $m$. It also adds $\phi(j)$ to the residual stream such that $j$ is the quotient $i/m$ if $i$ is a multiple of $m$. In the second layer, each position $i$ attends to the nearest position $j\leq i$ that is a multiple of $m$ and retrieves the (normalized) quotient stored there, which is $j/m=\lfloor i/m\rfloor$. It adds this (normalized) quotient in its own residual stream. We then use Lemma 7A.3) to construct a third layer that adds $\phi(i-1)$ and $\phi(i-2)$ to the residual stream. A fourth layer checks in parallel whether the quotient stored at $i$ matches the quotients stored at $i-1$ and $i-2$, respectively. In the fifth layer, position $i$ counts the number of positions storing the same quotient as $i$, excluding the first such position. Finally, in the sixth layer, position $i$ attends to position $a_{i}$ to compute and add to the residual stream $\phi(\lfloor a_{i}/m\rfloor)$ (which is $\phi(b_{i})$) and $\phi(a_{i}-m\lfloor a_{i}/m\rfloor)$ (which is $\phi(c_{i})$). We next describe a detailed implementation of the construction, followed by an argument of its correctness.
@@
Construction. The first layer uses an attention head with queries, keys, and values computed as follows. The query at position $i$ is $q_{i}=\phi(i,m)=\phi(i/m)$ computed via Lemma 3A.1) leveraging the assumption that $\psi(i)$ and $\psi(m)$ are present in the residual stream. The key and value at position $j$ are $k_{j}=v_{j}=\phi(j)$. Let $h^{1}_{i}=\phi(j)$ denote the head’s output. The layer adds $h^{1}_{i}$ to the residual stream and also adds $e_{i}=\mathbb{I}(h^{1}_{i}=\phi(i/m))$ using Lemma 8 (scalar equality check, §A.4) on the first coordinate of $h^{1}_{i}$ and $\phi(i/m)$. As we will argue in Appendix B, this layer has the intended behavior: $e_{i}=1$ if and only if $i$ is a multiple of $m$ and, if $e_{i}=1$, then the value it stores in the residual stream via $h^{1}_{i}$ is precisely the (normalized) quotient $\phi(i/m)$. [^4] As described in Lemma 8, a component will be added to the second layer to reset intermediate memory cells used in the first layer to 0 (this will happen analogously in later layers, but we will omit mentioning it).
@@
The second layer uses a head that attends with query $q_{i}=\langle 1,1\rangle$, key $k_{j}=\langle e_{j},[\phi(j)]_{0}\rangle$, and value $v_{j}=h^{1}_{j}$; note that both $e_{j}$ and $h^{1}_{j}$ can be read from the residual stream using masked pre-norm. This head attends to all positions $j\leq i$ that are multiples of $m$ (where $e_{j}=1$), with $[\phi(j)]_{0}$, the first component of $\phi(j)$, serving as a tie-breaking term for breaking ties in favor of the nearest multiple of $m$. As we will argue in Appendix B, this head outputs $\phi(\lfloor i/m\rfloor)$, which we store in the residual stream.
The third layer uses Lemma 7A.3) to add $\phi(i-1)$ and $\phi(i-2)$ to the residual stream at position $i$.
@@
In parallel for $k\in\{1,2\}$, the fourth layer attends with query $q_{i}=\phi(i-k)$, key $k_{j}=\phi(j)$, and value $v_{j}=\phi(\lfloor j/m\rfloor)$ to retrieve the quotient stored at position $i-k$. It uses Lemma 8A.4) on the first coordinate to store in the residual stream a boolean $b^{k}_{i}=\mathbb{I}(\phi(\lfloor i/m\rfloor)=\phi(\lfloor(i-k)/m\rfloor))$, indicating whether the quotient stored at $i$ matches the quotient stored at $i-k$.
@@
In the fifth layer, position $i$ attends with query $q_{i}=\langle\phi(\lfloor i/m\rfloor),1\rangle$, key $k_{j}=\langle\phi(\lfloor j/m\rfloor),b^{1}_{j}\rangle$, and value $v_{j}=1-b^{2}_{j}$. When the output of this layer is passed through layer norm, it produces $\phi(i\bmod m)$ as justified in Appendix B.
@@
The seventh and final layer cleans up any remaining intermediate values stored in the residual stream, setting them back to 0 as per Lemma 8. This is possible because all values $v$ are of the form $\phi(x)$ or a boolean, which means adding $-\phi(v)$ to the residual stream will reset the corresponding cell to 0.
Correctness. We justify the correctness of each layer in this construction in Appendix B. ∎
Our division construction is somewhat similar to the modular counting construction from Strobl et al. (2024), though the tools and underlying assumptions are different. Specifically, their approach relies on nonstandard position embeddings whereas ours uses masked pre-norm.

### v3

@@ 1 Introduction
A line of recent work analyzing the intrinsic computational power of transformers, the neural architecture behind today’s immensely successful large language models (LLMs), has established that, with fixed depth, transformers cannot express many simple problems outside the complexity class $\mathsf{TC}^{0}$, including recognizing regular languages and resolving connectivity of nodes in a graph (merrill-sabharwal-2023-parallelism; chiang2023tighter). These problems conceivably underlie many natural forms of sequential reasoning, such as state tracking (liu2023transformers; merrill2024illusion) and resolving logical inferences across long chains (wei2022chain). Thus, these results suggest inherent limitations on the types of reasoning transformer classifiers can perform. Yet, these findings come with an important caveat: even if transformers cannot solve such problems exactly for inputs of arbitrary lengths, they may still be able to solve them over inputs up to some bounded length. This perspective, coupled with the fact that treating depth as fixed is crucial to prior analyses placing transformers in $\mathsf{TC}^{0}$, motivates two related questions about depth as an important resource for a transformer, in relation to the context length over which it reasons:
@@
We address these questions by analyzing the expressive power of “universal” transformers (also called “looped” transformers) where a fixed model is given dynamic depth by repeating a block of middle layers a variable number of times (dehghani2018universal; yang2024looped; geiping2025depthscaling). We capture the regime where depth grows minimally with context length, with the middle layers repeated $\Theta(\log n)$ times on contexts of length $n$. We prove that even such highly uniform transformers, when allowed log-depth, can recognize regular languages [^1] While liu2023transformers provided a similar result, their construction relied heavily on an unconventional, non-uniform transformer architecture, requiring a different set of model weights for each input length. In contrast, our result holds in a stronger setting of highly uniform transformers—where model weights are not only fixed (independent of input length) as in practice but even shared across blocks of layers, enabling effective inference-time scaling. Our formal model incorporates standard architectural choices like residual connection and layer norm. We supplement our stronger theoretical findings with matching empirical learnability results. and solve graph connectivity, two important reasoning problems known to be beyond fixed-depth transformers (merrill-sabharwal-2023-parallelism). Our core technical contribution enabling this is Lemma˜1, that fully uniform transformers can compute division and remainder of small integers. This not only obviates the need for non-uniformity and special positional encodings relied upon in prior work, it is also an interesting finding on its own.
@@
Figure 1: To recognize a regular language over inputs of length $n$, the depth of a universal transformer can grow $\Theta(\log n)$ by Theorem˜1. On the other hand, width must grow superpolynomially (Theorem˜3), and the number of chain-of-thought steps must be superlogarithmic (Theorem˜4). The precise depth and width coefficients plotted here were obtained experimentally in Section˜7.
@@
Our result has two interesting interpretations. First, it directly shows that, by dynamically increasing their depth as $\Theta(\log n)$ on inputs of length $n$, one can construct transformers to solve regular language recognition (Theorem˜1) and graph connectivity (Theorem˜2) for arbitrary context length. In contrast, chain-of-thought (CoT) steps, used for additional test-time compute by newest LLMs such as OpenAI o1 (o1-short) and DeepSeek-R1 (deepseekr1-short), must be scaled superlogarithmically (Theorem˜4) to solve these problems, and width must be scaled superpolynomially (Theorem˜3), as shown in Figure˜1. Thus scaling depth more efficiently allows solving these reasoning problems compared to scaling width or using CoT.
@@
Second, a universal transformer unrolled to a fixed (independent of input size) depth $d$ is a special case of a standard $d$-depth transformer, namely one with a highly uniform structure (parameters shared across layers). Thus, our result shows that standard transformers with a fixed depth $d$ can recognize regular languages (Corollary˜1.2) and solve graph connectivity problems (Corollary˜2.1) as long as one cares only about bounded inputs of size $2^{O(d)}$. This allows us to quantify how many layers are necessary for a desired input size. For instance, Corollaries˜1.2 and 2.1 imply that with a depth of only $32$, such as in LLaMA 3.1 7B (llama3-short) and OLMo 7B (olmo-short), transformers can recognize regular languages up to strings of length $107$ and solve connectivity for graphs with up to $128$ vertices. A depth of $80$ (such as in LLaMA 3.1 70B) makes these input size limits practically unbounded: strings of length up to 440440K and graphs with up to 2.12.1B vertices. Empirically, our experiments supplementing our theoretical findings demonstrate that scaling depth as $\Theta(\log n)$ is necessary and sufficient for learning to recognize hard regular languages. [^2] Code: https://github.com/jopetty/word-problem/tree/willm/log-depth/log-depth.
@@ 2 Preliminaries: Universal Transformers
We consider $(s,r,t)$-universal transformers, which are defined to have $s$ fixed initial layers at the start, a sequence of $r$ layers that is repeated some number of times based on the input length, and a sequence of $t$ fixed final/terminal layers. Thus, an $(s,r,t)$-universal transformer unrolled $d(n)$ times for input length $n$ has a total of $s+rd(n)+t$ layers. geiping2025depthscaling empirically explored such transformers for scaling test-time computation for reasoning problems. A standard $d$-layer transformer is $(d,0,0)$-universal (equivalently, $(0,0,d)$-universal), while a standard universal transformer (dehghani2018universal; yang2024looped) is $(0,1,0)$-universal.
@@
    An embedding projection matrix $\mathbf{E}:\Sigma\to\mathbb{Q}^{m}$ and positional encoding function $\pi:\mathbb{N}\to\mathbb{Q}^{m}$, which we assume separates $1$ from other indices (merrill-sabharwal-2024-cot);
@@
We next define how the transformer maps a sequence $w_{1}\cdots w_{n}\in\Sigma^{n}$ to an output value $y\in\Sigma$; to do so, we will always specify that the transformer is unrolled to a specific depth function $d(n)$, which we will consider to be $d(n)=\lceil\log n\rceil$. [^3] Following computer science conventions, we let $\log n\triangleq\log_{2}n$. The computation is inductively defined by the residual stream $\mathbf{h}_{i}$: a cumulative sum of all layer outputs at each token $i$. In the base case, the residual stream $\mathbf{h}_{i}$ is initialized to $\mathbf{h}^{0}_{i}=\mathbf{E}(w_{i})+\pi(i)$. We then iteratively compute $s+rd(n)+t$ more layers, deciding which layer to use at each step as follows:
$L^{\ell}=\begin{cases}s\textrm{-layer}\;\ell&\textrm{if}\;1<\ell\leq s\\ r\textrm{-layer}\;((\ell-s-1)\bmod r)+1&\textrm{if}\;s<\ell\leq s+rd(n)\\ t\textrm{-layer}\;\ell-s-rd(n)&\textrm{otherwise.}\end{cases}$
@@
An $(s,r,t)$-transformer unrolled to some fixed depth can be viewed as a “uniform” special case of a fixed-depth transformer. Thus, constructions of dynamic-depth transformers (depth $d(n)$ for inputs of length $n$) imply that, given any bounded context length $N$, there also exists a fixed-depth transformer with depth $d(N)$ for the task at hand. The fact that this can be done with a looped transformer with dynamic depth is, in fact, a stronger condition that shows the construction is uniform, which is formally important as non-uniform models of computation can have very strong and unrealistic power (cf. merrill2022SatAttnTC0). In this way, our results about looped transformers will provide insights about standard, non-looped transformers with bounded context lengths.
@@
To make Definition˜1 well-defined, we will next describe the structure of the self-attention and feedforward sublayers that make up the structure of each transformer layer. Our definition of the transformer will have two minor differences from practice:
@@
    Masked pre-norm: We assume standard pre-norm (xiong2020on) but add a learned mask vector $\mathbf{m}\in\mathbb{R}^{m}$ that can select specific dimensions of the residual stream per sublayer.
Under masked pre-norm, the sublayer input will be read as sequence of normalized residual stream values $\mathbf{z}_{i}=\mathsf{layer\_norm}(\mathbf{m}\odot\mathbf{h}_{i})$, where $\odot$ is elementwise product and layer-norm can be standard layer-norm (ba2016layernormalization) or RMS norm (zhang2019rms). The sublayer then maps $\mathbf{z}_{1},\ldots,\mathbf{z}_{n}$ to a sequence $\delta_{1},\ldots,\delta_{n}$, and updates the residual stream as $\mathbf{h}^{\prime}_{i}=\mathbf{h}_{i}+\delta_{i}$.
@@
Positional Encodings.
We will assume no positional encodings (kazemnejad2023nope), but that there is a beginning of sequence (BoS) symbol at the start. As described by merrill-sabharwal-2024-cot, our constructions will generalize to any positional encoding as long as either BoS is provided or the first token’s positional encoding is linearly separable from all other positional encodings.
@@ 2.1 Memory Management in Universal Transformers
Appendix˜A discusses in detail how $\{-1,0,1\}$ values can be stored directly in the residual stream, while a general scalar $z$ can be stored either as $\psi(z)=\langle z,1,-z,-1\rangle$ in its unnormalized form or as the unit vector $\phi(z)=1/\sqrt{2}\cdot\psi(z)/\sqrt{z^{2}+1}$ in its normalized form (“layer-norm hash”; cf. merrill-sabharwal-2024-cot). Importantly, however $z$ is stored, when it is read using masked pre-norm, we obtain $\phi(z)$. In Appendix˜A, we show how numerical values represented using $\psi$ or $\phi$ can be easily written (Lemma˜4), read (Lemma˜2), and deleted (Lemmas˜5 and 6) from the residual stream. We will leverage these operations heavily in our theoretical constructions.
2.2 Numerical Datatype and Precision
Our constructions will involve working with scalars used as pointers to token positions, which will be stored in and retrieved from the residual stream (as discussed in Section˜2.1). We thus need $\Omega(\log n)$ bits of precision. We now formalize the underlying datatype we use for our constructions.
We assume scalars are encoded as strings in $\{0,1\}^{p}$, with $p=c\log n$ for some fixed $c>0$. Note that model parameters in our fully uniform setting cannot depend on $n$, but activations can. We assume there is some datatype $\mathbb{D}_{p}$ that assigns a numerical semantics for each string in $\{0,1\}^{p}$. For $x\in\mathbb{R}$, let $x]_{\mathbb{D}_{p}}$ be $x$ rounded into $\mathbb{D}_{p}$, i.e., the bitstring whose numerical value in $\mathbb{D}_{p}$ is closest to $x$ (breaking ties in favor of the higher value).
Our constructions will be agnostic to the underlying details of $\mathbb{D}_{p}$. Instead, we will minimally assume that, for some fixed $c$, all arithmetic operations in the transformer computation graph (addition, multiplication, division, $\exp$, and layer-norm) are $p$-precise for $p\geq c\log n$ in the following sense:
Definition 4 ($p$-Precise Operations).
Let $f:\mathbb{R}^{k}\to\mathbb{R}$ be an operation with $p$-precision realization $\tilde{f}:\mathbb{D}_{p}^{k}\to\mathbb{D}_{p}$. We say $\tilde{f}$ is $p$-precise if, for any $x_{1},\ldots,x_{k}\in\mathbb{R}$ exactly representable in $\mathbb{D}_{p}$,
$f(x_{1},\ldots,x_{k})]_{\mathbb{D}_{p}}=\tilde{f}([x_{1}]_{\mathbb{D}_{p}},\ldots,[x_{k}]_{\mathbb{D}_{p}}).$
To apply this definition, we view the summation in attention heads as an $n$-ary operation. We also view layer-norm as a single operation from $\mathbb{R}^{m}\to\mathbb{R}^{m}$.
Definition˜4 is naturally satisfied by the log-precision transformer model formalized by merrill2024illusion and used in earlier work (merrill-sabharwal-2023-parallelism; merrill2023logic), as long as sufficient precision is used internally to compute attention and layer-norm precisely. It is an open question how much precision is required by the internal primitive operations of summation and layer-norm to guarantee their output is $(c\log n)$-precise. We intentionally abstract away these low-level details here, especially because, in practice, additional precision is typically allocated in any case for attention and layer-norm computation (micikevicius2018mixed; olmo-short).
Finally, we briefly note how Definition˜4 relates to other datatype models. chiang2023tighter propose a polynomial-precision rational datatype, which satisfies Definition˜4 because the first $c\log n$ bits (and potentially more) are correct. In contrast, finite-precision transformers do not satisfy Definition˜4 because only $\mathrm{O}(1)$ bits are correct. In particular, while the mixed-precision model of yang2025kneedeep can precisely represent attention, it cannot store $(c\log n)$-bit values in the residual stream or as the output of layer-norm. It therefore does not satisfy Definition˜4 or suffice for our constructions.
@@ 3 Fixed Depth Transformers Can Divide Small Integers
Let $a_{i},b_{i},c_{i},m\in\mathbb{Z}^{\geq 0}$ be such that $a_{i}=b_{i}m+c_{i}$ where $a_{i}\leq i$ and $c_{i}<m$. Suppose $\psi(i)$, $\psi(m)$, and $\phi(a_{i})$ (or $\psi(a_{i})$) are present in the residual stream of a transformer at each token $i$. Then, there exists a block of 7 transformer layers with causally masked attention and masked pre-norm that, on any input sequence, adds $\phi(b_{i})$ and $\phi(c_{i})$ to the residual stream at each token $i$.
@@
The overall idea is as follows. In the first layer, each position $i$ outputs an indicator of whether it’s a multiple of $m$. It also adds $\phi(j)$ to the residual stream such that $j$ is the quotient $i/m$ if $i$ is a multiple of $m$. In the second layer, each position $i$ attends to the nearest position $j\leq i$ that is a multiple of $m$ and retrieves the (normalized) quotient stored there, which is $j/m=\lfloor i/m\rfloor$. It adds this (normalized) quotient in its own residual stream. We then use Lemma˜7A.3) to construct a third layer that adds $\phi(i-1)$ and $\phi(i-2)$ to the residual stream. A fourth layer checks in parallel whether the quotient stored at $i$ matches the quotients stored at i1i-1 and i2i-2, respectively. In the fifth layer, position $i$ counts the number of positions storing the same quotient as $i$, excluding the first such position. Finally, in the sixth layer, position $i$ attends to position $a_{i}$ to compute and add to the residual stream $\phi(\lfloor a_{i}/m\rfloor)$ (which is $\phi(b_{i})$) and $\phi(a_{i}-m\lfloor a_{i}/m\rfloor)$ (which is $\phi(c_{i})$). We next describe a detailed implementation of the construction, followed by an argument of its correctness.
@@
Construction. The first layer uses the following attention head. The query at position $i$ is $q_{i}=\phi(i,m)=\phi(i/m)$ computed via Lemma˜3A.1) leveraging the assumption that $\psi(i)$ and $\psi(m)$ are present in the residual stream. The key and value at position $j$ are $k_{j}=v_{j}=\phi(j)$ Let $h^{1}_{i}=\phi(j)$ denote the head’s output. The feedforward sublayer computes $e_{i}=\mathbb{I}(h^{1}_{i}=\phi(i/m))$ using Lemma˜8 (scalar equality check, §A.4) on the first coordinate of $h^{1}_{i}$ and $\phi(i/m)$. By Lemma˜9B), $e_{i}=1$ if and only if $i$ is a multiple of $m$ and, if $e_{i}=1$, then $h^{1}_{i}=\phi(i/m)$, i.e., it represents the quotient $i/m$. We store $h^{1}_{i}=\phi(i/m)$ and $e_{i}$ to the residual stream. [^4] As described in §8, a component will be added to the second layer to reset intermediate memory cells used in the first layer to 0 (this will happen analogously in later layers, but we will omit mentioning it).
@@
The second layer uses a head that attends with query $q_{i}=\langle 1,1\rangle$, key $k_{j}=\langle e_{j},[\phi(j)]_{0}\rangle$, and value $v_{j}=h^{1}_{j}$; both $e_{j}$ and $h^{1}_{j}$ can be read from the residual stream using masked pre-norm. This head attends to all positions $j\leq i$ that are multiples of $m$ (where $e_{j}=1$), with $\phi(j)]_{0}$, the first component of $\phi(j)$, serving as a tie-breaking term for breaking ties in favor of the nearest multiple of $m$. Lemma˜10B) shows this head outputs $h^{2}_{i}=\phi(\lfloor i/m\rfloor)$, which we store in the residual stream.
The third layer uses Lemma˜7A.3) to add $\phi(i-1)$ and $\phi(i-2)$ to the residual stream at position $i$.
@@
In parallel for $k\in\{1,2\}$, the fourth layer attends with query $q_{i}=\phi(i-k)$, key $k_{j}=\phi(j)$, and value $v_{j}=\phi(\lfloor j/m\rfloor)$ to retrieve the quotient stored at position iki-k. It uses Lemma˜8 (SA.4) on the first coordinate to store in the residual stream a boolean $b^{k}_{i}=\mathbb{I}(\phi(\lfloor i/m\rfloor)=\phi(\lfloor(i-k)/m\rfloor))$, indicating whether the quotient stored at $i$ matches the quotient stored at iki-k.
@@
In the fifth layer, position $i$ attends with query $q_{i}=\langle\phi(\lfloor i/m\rfloor),1\rangle$, key $k_{j}=\langle\phi(\lfloor j/m\rfloor),b^{1}_{j}\rangle$, and value $v_{j}=1-b^{2}_{j}$. When the output $h^{5}_{i}$ of this layer is read through layer norm, it produces $\phi(h^{5}_{i})=\phi(i\bmod m)$ as proved in Lemma˜11B).
@@
The seventh and final layer cleans up any remaining intermediate values stored in the residual stream, setting them back to 0 as per Lemma˜8. This is possible because all values $v$ are of the form $\phi(x)$ or a boolean, which means adding $\phi(v)$ to the residual stream will reset the corresponding cell to 0. ∎
Our division construction is somewhat similar to the modular counting construction from strobl2024transformers, though the tools and underlying assumptions are different. Specifically, their approach relies on nonstandard position embeddings whereas ours uses masked pre-norm.